OpenStudy (anonymous):

Find the largest factor of 2520 that is not divisible by 6.

OpenStudy (anonymous):

315

OpenStudy (anonymous):

how do u calculate it ?

OpenStudy (alexwee123):

well the prime factorization of 2520 is 2^3 x 3^2 x 5 x 7 any factor of 2520 that is divisible by 6 contains at least one 2 and at least one 3 in its prime factorization. so remove factors so that no 2's remain or no 3's remain if we remove all the 3s, we divide the number by 3^2 = 9. but if we remove all the 2s, we divide the number by 2^3 = 8 since 8 is less than 9 the last factor 2520/8 = 315

OpenStudy (anonymous):

How did i calculate .... 2 x 2 x2 x 3 x 3 x5 x 7 the remainder value at first occurance of 3 is 315 which is not divisible by 6 first 3 occurance of 2 ... remainder values are divisible by 6 does it make sense